OpenStudy (anonymous):

the equation of a line that passes through (-2,2) and is perpendicular to y= -1/2 +5

OpenStudy (anonymous):

Hiya Peadchcat. Problem: the equation of a line that passes through (-2,2) and is perpendicular to y= -1/2 +5 I would approach this by breaking it down into steps. Two lines that are perpendicular, hmm. When their slopes are multiplied, you get -1. So the line that is given has a slope of -1/2. So the slope of the unknown line will be 2. OK, so the equation so far will be: y = 2x + b And "b" would stand for the y-intercept. So now I can plug in the x- and y-values known to fall on the second line: 2 = 2(-2) + b Then I would solve. 2 = -4 + b b = 2 So the equation is y = 2x + b. And I can check: yup, it has a slope of 2, so it'll be perpendicular to the first line, and the point (-2,2) makes the equation true. Done! Hope that helps!

OpenStudy (anonymous):

I meant "Peachcat"! :-)